# Learnable In-Context Vector for Visual Question Answering

### v1

@@ 4.3 Ablation Studies
Training Loss: Table 2 compares the results of using different losses: only $\mathcal{L}_{gt}$ in Eq. (7) or $\mathcal{L}_{d}$ in Eq. (7). We find that only using $\mathcal{L}_{gt}$ (same as standard fine-tuning) significantly damages the performance, e.g., $\mathcal{L}_{gt}$ achieves 16.9/6.12 lower accuracy on VQAv2/OKVQA
Table 2: Accuracy (%) of L-ICV with Different Training Loss on VQA.
@@
compared to using the combined loss $\mathcal{L}$; yet using only $\mathcal{L}_{d}$ results in a smaller performance drop – 3.78/3.14 lower VQAv2/OKVQA than when using $\mathcal{L}$. This suggests that with a limited number of trainable parameters, L-ICV trained with $\mathcal{L}_{d}$ is more robust and capable of capturing essential task information than $\mathcal{L}_{gt}$. It underscores that the L-ICV cannot solely rely on fine-tuning with LMMs on specific datasets, but should effectively leverage abstracted insights from demonstrations.
Number of Demonstrations $k$: We compare the performance of L-ICV trained with $k$ demonstrations per query and the corresponding $k$-shot ICL in Table 3. The result shows that an increase
@@
in the number of demonstrations enhances the performance of ICL and L-ICV, indicating that more demonstrations can provide each query with a richer context to help train L-ICV. Additionally, L-ICV consistently surpasses the performance of $k$-shot ICL across different training sizes, showcasing the robustness of our L-ICV in utilizing demonstrations. Notably, when the number of demonstrations is limited, the performance gap between L-ICV and ICL becomes more pronounced. This is because ICL is highly sensitive to the choice of demonstrations; with insufficient demonstrations, the model may shift the query representations in an incorrect direction. In contrast, L-ICV continuously by learning the main shift direction of the query representations from the demonstrations, reduces the negative impact of poor demonstrations on the query during training and is more robust that can extract essential task information.
Size of Training Set: Figure 4 illustrates how varying the number of training samples impacts the performance of L-ICV and LoRA. On the VQAv2 dataset, both methods show improved performance with increasing data sizes. Notably, L-ICV performs exceptionally well across both low and high training sizes. It achieves performance close to that of 1-shot ICL with just 700 training samples and surpasses 32-shot ICL with 4,000 training samples. In contrast, LoRA does not exceed the performance of 1-shot ICL, even when expanded to 8,000 samples. For OKVQA, the performance of LoRA with small data sizes is even worse than Zero-Shot. This is because OKVQA requires external knowledge to answer questions, while learning external knowledge from a small amount of data can disrupt the inherent knowledge of the pre-trained model, leading to a significant drop in performance. Conversely, L-ICV excels by focusing on learning shift direction, thus preserving the model’s inherent reasoning abilities. With just 500 samples, L-ICV outperforms 1-shot ICL, and with 4,000 samples, it nearly matches the performance of 32-shot ICL. These observations underscore L-ICV’s superior efficiency over LoRA in capturing and utilizing complex reasoning capabilities with much fewer training samples.
@@ 4.4 Analysis
Figure 4: Accuracy (%) of L-ICV and LoRA with different size of training set.
@@ 4.4.1 The Shifting Effect in Latent Space
To better demonstrate the shifting effect of L-ICV on query samples, we randomly select 200 query samples and conduct different methods of inference in LMMs. We extract the representation vector of the first answer token for T-SNE dimensionality reduction, shown in Figure 5. Additionally, to quantitatively evaluate the effect of shift directions of different ICV methods, we calculate the following metrics. Given a query $\hat{\bm{x}}$, we use $\bm{r}_{icl}$, $\bm{r}_{zs}$, $\bm{r}^{*}$ to denote the representation of the first answer token obtained by 32-shot ICL, Zero-Shot, specific ICV methods, respectively. Then we calculate the standard shift direction as $\bm{s}_{gt}=\bm{r}_{icl}-\bm{r}_{zs}$ and the shift direction of specific ICV as $\bm{s}^{*}=\bm{r}^{*}-\bm{r}_{zs}$. Finally, we define the shift direction similarity as the cosine similarity between $\bm{s}^{*}$ and $\bm{s}_{gt}$, indicating how closely the shift direction of the ICV method aligns with the standard shift direction. The results are presented in Table 4.
@@
From Figure 5, we can find that 32-shot ICL exhibits a significant shift compared to Zero-Shot, visualizing the shift effects given in Eq. 4. Considering both Table 4 and 1, we find that the shift direction similarity has positive correlation to the accuracy: if a method has large direction similarity, it also has better performance. For example, among non-learnable methods, TV has higher shift direction similarity than other ones, then it has better accuracy in Table 1. Furthermore, for L-ICV which has the best accuracy in Table 1, its shift direction similarity is also the highest, which is 0.742/0.829 on VQAv2/OKVQA, validating that L-ICV can produce shifts in query samples similar to 32-shot ICL, as visualized in Figure 5. Such positive correlation validates the effectiveness of our motivation that a single L-ICV can indeed simulate the ICL capability of LMMs by shifting the direction of the query representation.
@@ 4.4.2 Why Non-Learnable Methods are Poor on VQA?
L-ICV ’Question’, ’_Short’, ’?’, ’no’, ’QUEST’, ’questions’, ’\$?’, ’answer’, ’Short’, ’_questions’
@@
We calculate the $\bm{p}$ of the vectors got from different methods in VQAv2 and select the top-$10$ tokens with the highest probabilities in $\bm{p}$ shown in Table 5. We can see that the tokens got from FV are not highly relevant to the VQA task, which proves that FV does not capture the task information of VQAv2. On the other hand, the frequency of "yes" and "no" tokens is relatively high in the decoding results of PCA-ICV and TV, suggesting that they prefer to capture the simple patterns from demonstrations, e.g.yes/no, but struggle to grasp the overall task information of complex VQA. In contrast, the tokens of L-ICV decoding are not biased to specific answers like yes/no, suggesting it abstracts more summary task knowledge of VQA.
@@ 5 Conclusion
To address the two major drawbacks of ICL in LMM—long computation time and sensitivity to demonstration selection—we try to apply non-learnable ICV methods from NLP to solve VQA. However, due to the complexity of VQA and the significant biases often inherent in non-learnable methods, the performance is unsatisfactory. Then we propose the Learnable ICV (L-ICV) to overcome this drawback. By learning the general shift direction from a large amount of ICL data, L-ICV successfully replaces the role of demonstrations in ICL. Experiments validate that L-ICV outperforms traditional ICL methods and other non-learnable ICV methods on two VQA datasets. Experiments also show that L-ICV, compared to LoRA, maintains excellent performance with minimal data, suggesting L-ICV is a new research direction for LMMs to solve multimodal tasks. In the future, we will explore the application of L-ICV on more multimodal tasks by various LMMs.
@@
A.1 L-ICV Hyperparameters
Table 7 presents the hyper-parameters utilized for training the L-ICV. The optimizer denotes the optimization algorithm employed during model training. For $\bm{V}$ and $\bm{\alpha}$, we use different learning rate to optimize. The $\lambda$ represents the weight assigned to $\mathcal{L}_{\text{gt}}$ in Eq 7 during training. Precision refers to the float precision type used for model weights and gradient descent throughout the ICV training process. Weight Decay signifies the rate of weight decay applied during training, the warm up value is set to 0.01. while accumulate batches denotes the batch size for gradient accumulation during the training phase.
Table 7: VQAv2 and OKVQA L-ICV Training Parameters
@@ B.1 The Detailed Inference Speed Experiments
This subsection provides a detailed presentation of some experimental data, primarily including forward propagation FLOPs and runtime cost, as well as a comparison of specific data between L-ICV and LoRA. Table 9 presents the detailed results of the runtime and the FLOPs shown in Figure 3. For a forward operation L-ICV compared to k-shot ICL. The average token length during the forward inference is 38 tokens for Zero-Shot and L-ICV, 107 tokens for 4-shot ICL, 187 tokens for 8-shot ICL, 330 tokens for 16-shot ICL, and 633 tokens for 32-shot ICL.
Table 9: Comparison of FLOPs and Runtime for L-ICV and $k$-shot ICL
Mertic Zero-Shot L-ICV 4-shot ICL 8-shot ICL 16-shot ICL 32-shot ICL
@@ B.2 The Detailed Accuracy of Different Training Dataset
Table 10 presents a comparative analysis of the results between LoRA and L-ICV results on different training dataset size in Section 4.3.
@@ B.3 Non-Learnable Methods Results
Table 10: Accuracy (%) of LoRA and L-ICV on different training set sizes
@@
Appendix C General L-ICV
L-ICV is essentially a shift vector, allowing us to control the shift directions of query representations, which means we can control the shift direction of LMMs by adding or subtracting different shift vectors. Therefore, we can use several L-ICV trained on different VQA dataset to get the General L-ICV. Specifically, given the diverse VQA task set $\bm{T}=\{\bm{t}_{1},\dots,\bm{t}_{n}\}$, where $\bm{t}_{i}$ is a VQA task and $n$ is the number of tasks, we train the task-specific L-ICV $\bm{V}^{i}=\{\bm{v}_{1}^{i},\dots,\bm{v}_{L}^{i}\}$ and the weight factor $\bm{\alpha}^{i}=\{\alpha_{1}^{i},\dots,\alpha_{L}^{i}\}$ on each VQA dataset. Then, we have the L-ICV set $\mathcal{V}_{set}=\{\bm{V}^{1},\dots,\bm{V}^{n}\}$ and its weight factors set $\bm{\alpha}_{set}=\{\bm{\alpha}^{1},\dots,\bm{\alpha}^{n}\}$. The General L-ICV has the same shape of task-specific L-ICV. For the $l$-th layer, it is defined as $\bm{v}_{l}=\sum_{i}\bm{\alpha}_{l}^{i}\bm{v}_{l}^{i}$. During inference, we use the vector $\bm{v}_{l}$ to shift the original representation, resulting in $h(\hat{x}_{i})^{\prime}=h(\hat{x}_{i})+\bm{v}_{l}$. We average the L-ICV trained in OKVQA and VQAv2 to get the general ICV and evaluate the performance of the general ICV $\bm{V}_{g}=\{\bm{v}_{1},\dots,\bm{v}_{L}\}$ on OKVQA and VQAv2, with the results presented on Table 16.
Table 16: Accuracy (%) for 32-shot ICL, task-specific L-ICV, General L-ICV.
Methods 32-shot ICL L-ICV General L-ICV
@@ Appendix C General L-ICV
The results indicate that while the performance of the general ICV is somewhat reduced compared to task-specific L-ICV; it decreased by 2.37 on VQAv2 and by 0.56 on OKVQA. However, its performance is very close to that of 32-shot ICL. Most importantly, it offers significant advantages in real-world scenarios where the distribution of test data is unknown. This makes the general ICV more suitable for practical applications, providing a robust solution that can be effectively utilized across varying environments. Furthermore, the findings highlight the scalability of L-ICV: whenever a new VQA dataset is used to train L-ICV, we only need to simply recalculate the mean shift vectors of all previously trained task-specific L-ICVs and the new L-ICV to generate a new general ICV, thus creating a more general VQA L-ICV. This approach not only simplifies the adaptation process for diverse tasks but also ensures that the model maintains a high level of performance across different applications. The results, therefore, underscore the potential of general ICVs to enhance the flexibility and applicability of VQA systems in real-world settings.

### v3

@@ 4.3 Ablation Studies
Training Loss: Table 2 compares the results of using different losses: only $\mathcal{L}_{gt}$ in Eq. (7) or $\mathcal{L}_{d}$ in Eq. (7). We find that only using $\mathcal{L}_{gt}$ (same as standard fine-tuning) significantly damages the performance, e.g., $\mathcal{L}_{gt}$ achieves 16.9/6.12 lower accuracy on VQAv2/OKVQA compared to using the combined loss $\mathcal{L}$; yet using only $\mathcal{L}_{d}$ results in a smaller performance drop – 3.78/3.14 lower VQAv2/OKVQA than when using $\mathcal{L}$. This suggests that with a limited number of trainable parameters, LIVE trained with $\mathcal{L}_{d}$ is more robust and capable of capturing essential task information than $\mathcal{L}_{gt}$. It underscores that the LIVE cannot solely rely on fine-tuning with LMMs on specific datasets, but should effectively leverage abstracted insights from demonstrations.
Table 2: Accuracy (%) of LIVE with Different Training Loss on VQA.
@@
Number of Demonstrations $k$: We compare the performance of LIVE trained with $k$ demonstrations per query and the corresponding $k$-shot ICL in Table 3. The result shows that an increase
@@
in the number of demonstrations enhances the performance of ICL and LIVE, indicating that more demonstrations can provide each query with a richer context to help train LIVE. Additionally, LIVE consistently surpasses the performance of $k$-shot ICL across different training sizes, showcasing the robustness of our LIVE in utilizing demonstrations. Notably, when the number of demonstrations is limited, the performance gap between LIVE and ICL becomes more pronounced. This is because ICL is highly sensitive to the choice of demonstrations; with insufficient demonstrations, the model may shift the query representations in an incorrect direction. In contrast, LIVE continuously by learning the main shift direction of the query representations from the demonstrations, reduces the negative impact of poor demonstrations on the query during training and is more robust that can extract essential task information.
Size of Training Set: Figure 4 illustrates how varying the number of training samples impacts the performance of LIVE and LoRA. On the VQAv2 dataset, both methods show improved performance with increasing data sizes. Notably, LIVE performs exceptionally well across both low and high training sizes. It achieves performance close to that of 1-shot ICL with just 700 training samples and surpasses 32-shot ICL with 4,000 training samples. In contrast, LoRA does not exceed the performance of 1-shot ICL, even when expanded to 8,000 samples. For OKVQA, the performance of LoRA with small data sizes is even worse than Zero-Shot. This is because OKVQA requires external knowledge to answer questions, while learning external knowledge from a small amount of data can disrupt the inherent knowledge of the pre-trained model, leading to a significant drop in performance. Conversely, LIVE excels by focusing on learning shift direction, thus preserving the model’s inherent reasoning abilities. With just 500 samples, LIVE outperforms 1-shot ICL, and with 4,000 samples, it nearly matches the performance of 32-shot ICL. These observations underscore LIVE’s superior efficiency over LoRA in capturing and utilizing complex reasoning capabilities with much fewer training samples.
@@ 4.4 Analysis
Figure 4: Accuracy (%) of LIVE and LoRA with different size of training set.
@@ 4.4.1 The Shifting Effect in Latent Space
To better demonstrate the shifting effect of LIVE on query samples, we randomly select 200 query samples and conduct different methods of inference in LMMs. We extract the representation vector of the first answer token for T-SNE dimensionality reduction, shown in Figure 5. Additionally, to quantitatively evaluate the effect of shift directions of different ICV methods, we calculate the following metrics. Given a query $\hat{\bm{x}}$, we use $\bm{r}_{icl}$, $\bm{r}_{zs}$, $\bm{r}^{*}$ to denote the representation of the first answer token obtained by 32-shot ICL, Zero-Shot, specific ICV methods, respectively. Then we calculate the standard shift direction as $\bm{s}_{gt}=\bm{r}_{icl}-\bm{r}_{zs}$ and the shift direction of specific ICV as $\bm{s}^{*}=\bm{r}^{*}-\bm{r}_{zs}$. Finally, we define the shift direction similarity as the cosine similarity between $\bm{s}^{*}$ and $\bm{s}_{gt}$, indicating how closely the shift direction of the ICV method aligns with the standard shift direction. The results are presented in Table 4.
@@
From Figure 5, we can find that 32-shot ICL exhibits a significant shift compared to Zero-Shot, visualizing the shift effects given in Eq. 4. Considering both Table 4 and 1, we find that the shift direction similarity has positive correlation to the accuracy: if a method has large direction similarity, it also has better performance. For example, among non-learnable methods, TV has higher shift direction similarity than other ones, then it has better accuracy in Table 1. Furthermore, for LIVE which has the best accuracy in Table 1, its shift direction similarity is also the highest, which is 0.742/0.829 on VQAv2/OKVQA, validating that LIVE can produce shifts in query samples similar to 32-shot ICL, as visualized in Figure 5. Such positive correlation validates the effectiveness of our motivation that a single LIVE can indeed simulate the ICL capability of LMMs by shifting the direction of the query representation.
@@ 4.4.2 Why Non-Learnable Methods are Poor on VQA?
LIVE ’Question’, ’_Short’, ’?’, ’no’, ’QUEST’, ’questions’, ’\$?’, ’answer’, ’Short’, ’_questions’
@@
We calculate the $\bm{p}$ of the vectors got from different methods in VQAv2 and select the top-$10$ tokens with the highest probabilities in $\bm{p}$ shown in Table 5. We can see that the tokens got from FV are not highly relevant to the VQA task, which proves that FV does not capture the task information of VQAv2. On the other hand, the frequency of "yes" and "no" tokens is relatively high in the decoding results of PCA-ICV and TV, suggesting that they prefer to capture the simple patterns from demonstrations, e.g.yes/no, but struggle to grasp the overall task information of complex VQA. In contrast, the tokens of LIVE decoding are not biased to specific answers like yes/no, suggesting it abstracts more summary task knowledge of VQA.
@@ 5 Conclusion
To address the two major drawbacks of ICL in LMM—long computation time and sensitivity to demonstration selection—we try to apply non-learnable ICV methods from NLP to solve VQA. However, due to the complexity of VQA and the significant biases often inherent in non-learnable methods, the performance is unsatisfactory. Then we propose the Learnable ICV (LIVE) to overcome this drawback. By learning the general shift direction from a large amount of ICL data, LIVE successfully replaces the role of demonstrations in ICL. Experiments validate that LIVE outperforms traditional ICL methods and other non-learnable ICV methods on two VQA datasets. Experiments also show that LIVE, compared to LoRA, maintains excellent performance with minimal data, suggesting LIVE is a new research direction for LMMs to solve multimodal tasks. In the future, we will explore the application of LIVE on more multimodal tasks by various LMMs.
Acknowledgement
This work is supported by National Science Foundation of China (62206048), Natural Science Foundation of Jiangsu Province (BK20220819), Young Elite Scientists Sponsorship Program of Jiangsu Association for Science and Technology Tj-2022-027, Fundamental Research Funds for the Central Universities(2242024k30035) and Big Data Computing Center of Southeast University
@@ Acknowledgement
A.1 LIVE Hyperparameters
Table 7 presents the hyper-parameters utilized for training the LIVE. The optimizer denotes the optimization algorithm employed during model training. For $\bm{V}$ and $\bm{\alpha}$, we use different learning rate to optimize. The $\lambda$ represents the weight assigned to $\mathcal{L}_{\text{gt}}$ in Eq 7 during training. Precision refers to the float precision type used for model weights and gradient descent throughout the ICV training process. Weight Decay signifies the rate of weight decay applied during training, the warm up value is set to 0.01. while accumulate batches denotes the batch size for gradient accumulation during the training phase.
Table 7: VQAv2 and OKVQA LIVE Training Parameters
@@ B.1 The Detailed Inference Speed Experiments
This subsection provides a detailed presentation of some experimental data, primarily including forward propagation FLOPs and runtime cost, as well as a comparison of specific data between LIVE and LoRA. Table 9 presents the detailed results of the runtime and the FLOPs shown in Figure 3. For a forward operation LIVE compared to k-shot ICL. The average token length during the forward inference is 38 tokens for Zero-Shot and LIVE, 107 tokens for 4-shot ICL, 187 tokens for 8-shot ICL, 330 tokens for 16-shot ICL, and 633 tokens for 32-shot ICL.
Table 9: Comparison of FLOPs and Runtime for LIVE and $k$-shot ICL
Mertic Zero-Shot LIVE 4-shot ICL 8-shot ICL 16-shot ICL 32-shot ICL
@@ B.2 The Detailed Accuracy of Different Training Dataset
Table 10 presents a comparative analysis of the results between LoRA and LIVE results on different training dataset size in Section 4.3.
@@ B.3 Non-Learnable Methods Results
Table 10: Accuracy (%) of LoRA and LIVE on different training set sizes
@@
Appendix C General LIVE
LIVE is essentially a shift vector, allowing us to control the shift directions of query representations, which means we can control the shift direction of LMMs by adding or subtracting different shift vectors. Therefore, we can use several LIVE trained on different VQA dataset to get the General LIVE. Specifically, given the diverse VQA task set $\bm{T}=\{\bm{t}_{1},\dots,\bm{t}_{n}\}$, where $\bm{t}_{i}$ is a VQA task and $n$ is the number of tasks, we train the task-specific LIVE $\bm{V}^{i}=\{\bm{v}_{1}^{i},\dots,\bm{v}_{L}^{i}\}$ and the weight factor $\bm{\alpha}^{i}=\{\alpha_{1}^{i},\dots,\alpha_{L}^{i}\}$ on each VQA dataset. Then, we have the LIVE set $\mathcal{V}_{set}=\{\bm{V}^{1},\dots,\bm{V}^{n}\}$ and its weight factors set $\bm{\alpha}_{set}=\{\bm{\alpha}^{1},\dots,\bm{\alpha}^{n}\}$. The General LIVE has the same shape of task-specific LIVE. For the $l$-th layer, it is defined as $\bm{v}_{l}=\sum_{i}\bm{\alpha}_{l}^{i}\bm{v}_{l}^{i}$. During inference, we use the vector $\bm{v}_{l}$ to shift the original representation, resulting in $\mathcal{M}_{l}(\hat{x}_{i})^{\prime}=\mathcal{M}_{l}(\hat{x}_{i})+\bm{v}_{l}$. We average the LIVE trained in OKVQA and VQAv2 to get the general ICV and evaluate the performance of the general ICV $\bm{V}_{g}=\{\bm{v}_{1},\dots,\bm{v}_{L}\}$ on OKVQA and VQAv2, with the results presented on Table 16.
Table 16: Accuracy (%) for 32-shot ICL, task-specific LIVE, General LIVE.
Methods 32-shot ICL LIVE General LIVE
@@ Appendix C General LIVE
The results indicate that while the performance of the general ICV is somewhat reduced compared to task-specific LIVE; it decreased by 2.37 on VQAv2 and by 0.56 on OKVQA. However, its performance is very close to that of 32-shot ICL. Most importantly, it offers significant advantages in real-world scenarios where the distribution of test data is unknown. This makes the general ICV more suitable for practical applications, providing a robust solution that can be effectively utilized across varying environments. Furthermore, the findings highlight the scalability of LIVE: whenever a new VQA dataset is used to train LIVE, we only need to simply recalculate the mean shift vectors of all previously trained task-specific LIVEs and the new LIVE to generate a new general ICV, thus creating a more general VQA LIVE. This approach not only simplifies the adaptation process for diverse tasks but also ensures that the model maintains a high level of performance across different applications. The results, therefore, underscore the potential of general ICVs to enhance the flexibility and applicability of VQA systems in real-world settings.
Appendix D More exploratory experiments.
D.1 Generalization Ability
We conducted supplementary experiments on other vision-language tasks and LMMs to demonstrate the general applicability of our LIVE method.
D.1.1 Task Generalization
In Table 17, we utilized IDEFICSv1-9B as the baseline model to evaluate the performance of LIVE against ICL and LoRA on the COCO caption dataset[54], OcrVQA dataset[55], and VL-ICL dataset[56]. Experimental results demonstrate that our method consistently maintains advantages over both ICL and LoRA across various tasks.
Table 17: Accuracy (%) for 32-shot ICL, LoRA, and LIVE in various dataset.
Task (Metric) 32-shot ICL LoRA LIVE
COCO (CIDEr) 106.31 109.18 117.38
VL-ICL Textocr (ACC) 21.5 22.5 24.0
VL-ICL Clevr (ACC) 33.5 35.5 37.0
OcrVQA (ACC) 16.3 15.9 17.5
D.1.2 Model Generalization
To test the generalizability of our method across different architectural models, we extended LIVE to IDEFICSv2 model[57]. It is noteworthy that IDEFICSv2 (8B-base) differs from IDEFICSv1 in its handling of image-text interactions; while IDEFICSv1 uses a cross-attention mechanism for the fusion of image and text tokens, IDEFICSv2 utilizes an additional projection layer to map image information into the text space for LMM processing. These two architectures represent the primary approaches within LMM. Conducting experiments with this model can validate the generalizability of LIVE.
Due to the longer input sequence of IDEFICSv2 compared to IDEFICSv1, we reached the limits of our GPU memory during 8-shot ICL inference. Therefore, we compared the LIVE trained on 8-shot data with 1-8 shot ICL. As shown in Table  18, LIVE improves upon 8-shot ICL by 6.1% and 0.84% on VQAv2 and OKVQA, respectively. In terms of inference speed, it surpasses 8-shot ICL by a factor of 8.95. These results demonstrate that the LIVE method is both generalizable and effective within LMM, excelling in both inference accuracy and speed compared to ICL.
Table 18: The performance comparison between LIVE and ICL on IDEFICS-v2-8B.
Method VQAv2 ACC OKVQA ACC Inference Time FLOPs (TFLOPs)
Zero-shot 55.39 43.08 0.087 ($\times$ 0.97) 3.110 ($\times$ 0.99)
1-shot ICL 60.33 45.65 0.168 ($\times$ 1.90) 6.497 ($\times$ 2.07)
2-shot ICL 63.49 50.40 0.237 ($\times$ 2.68) 9.610 ($\times$ 3.06)
4-shot ICL 64.88 53.18 0.395 ($\times$ 4.46) 17.139 ($\times$ 5.46)
8-shot ICL 66.20 57.68 0.792 ($\times$ 8.95) 32.130 ($\times$ 10.23)
8-shot LIVE 70.30 58.52 0.089 ($\times$ 1) 3.141 ($\times$ 1)
D.2 Performance with more shot
To investigate the impact of a higher number of shots on LIVE, we conducted a comparison with ICL at 48 and 64 shots. Detail results are shown in Table 19 .The results indicate that while increasing the number of shots improves the performance of both methods, they have reached a performance plateau. Notably, the gains achieved by LIVE remain superior to those of ICL. For example, with 64 shots LIVE achieves a 2.99%/1.85% improvement over ICL on VQAv2/OKVQ. Thus, although using more ICDs improves ICL performance, LIVE has a large improvement
Table 19: The performance comparison of ICL and LIVE with more shots.
Benchmark Method Shot 16 Shot 32 Shot 48 Shot 64
VQAv2 ICL 55.7 56.18 56.67 56.71
LIVE 58.27 (+2.57) 58.54 (+2.36) 59.07 (+2.4) 59.70 (+2.99)
OKVQA ICL 47.7 48.48 48.68 48.60
LIVE 49.71 (+2.01) 50.08 (+1.6) 50.55 (+1.87) 50.45 (+1.85)
D.3 Shared layer LIVE
We conducted experiments to investigate the impact of using a shared bias vector across all layers (as shown in Table 20), and we observed a significant performance drop compared to LIVE: 20.6% on VQAv2 and 9.01% on OKVQA. This performance drop aligns with the findings in [24] that different Transformer layers play diverse roles in ICL. Also, Eq 5 show that LIVE vectors are layer-specific. Then using a single shared vector fails to capture the distinct information processed at each layer, leading to the observed decline in accuracy.
Table 20: The performance comparison of using a shared LIVE across all layer (Share LIVE) with LIVE and ICL
Method VQAv2 OKVQA
32-shot ICL 56.18 48.48
32-shot LIVE 58.54 50.08
32-shot Share LIVE 37.94 (-20.6) 41.07 (-9.01)
D.4 Detail comparisopn of LoRA
To provide a more detailed comparison with LoRA, we made the following modifications to LoRA. First, based on the loss function formula in Eq. 7 we applied the same loss function to LoRA. As shown in Table 21, even after modifying the loss function, LoRA’s performance remains significantly lower than LIVE on both VQAv2 (49.48% vs. 58.54%) and OKVQA (37.02% vs. 50.08%). Additionally, we fine-tuned LoRA with more parameters, but similarly, increasing the parameter count did not close the performance gap between LoRA and our LIVE, further demonstrating the effectiveness of LIVE’s unique design.
Table 21: The performance comparison of the different implementations of LoRA with LIVE. “LoRA” integrates LoRA into the token classification head of the last layer to minimize the trainable parameters, “LoRA with KL” uses 32-shot ICDs and applies the same KL loss as our LIVE, and “LoRA O_proj all layers” integrates LoRA into the O_proj of all layers.
Method VQAv2 OKVQA Parameters Number
LoRA 49.02 34.21 1.15M (x 8 times)
LoRA with KL 49.48 37.02 1.15M (x 8 times)
LoRA O_proj all layers 57.57 45.03 10.5M (x 80 times)
LIVE 58.54 50.08 0.13M (x 1 time)
D.5 Comparison between untrained LIVE
We evaluated the impact of untrained LIVE by initializing them randomly with a normal distribution and tested on VQAv2 (Table 22), revealing that untrained LIVE has similar results as zero-shot, which are largely worse than the trained LIVE. We also visualized the embeddings using t-SNE to see the shift effect in Fig 7 (A), showing that the shift effect of untrained LIVE is similar to zero-shot ICL, while learned LIVEs align more closely with 32-shot ICL, proving that the training process is important.
Table 22: The performance comparison of untrained LIVE, trained LIVE, ICL, and Zero-shot.
Task Zero Shot Untrained LIVE 32-shot ICL Trained LIVE (Paper)
VQAv2 29.25 30.3 56.18 58.54
OKVQA 30.54 30.2 48.48 50.08
Figure 7: T-SNE visualization of first answer token over 200 queries. Left/right corresponds to VQAv2/OKVQA, respectively.
Table 23: The performance comparison between LIVE and diverse ICL methods.
Task Method Shot 1 Shot 4 Shot 8 Shot 16 Shot 32 Average
VQAv2 LIVE 56.84 57.60 58.25 58.27 58.54 57.90
ICL 51.39 53.72 54.24 55.70 56.18 54.25
RICE 47.83 53.54 55.04 56.89 58.07 54.27
MMICES 48.22 54.99 56.16 57.02 57.98 54.87
OKVQA LIVE 47.51 47.68 49.4 49.71 50.08 48.88
ICL 40.75 46.11 46.79 47.70 48.48 45.97
RICE 40.67 47.07 48.92 50.73 51.11 47.70
MMICES 40.39 47.13 50.19 50.46 50.61 47.76
D.6 Comparison between LIVE and diverse ICL methods.
In previous works[8, 19, 18], the importance of selecting ICDs for ICL in VQA and other vision-language tasks has been demonstrated. To reasonably compare the performance differences between LIVE and ICL, we adopted two methods for selecting ICL samples: RICE[43], which retrieves samples based on image similarity to the query, and MMICES[58], which combines both language and image similarity for ICD retrieval. The results comparing LIVE and ICL with different ICD selection strategies are presented in the Table 23, showing that even with optimized ICD selection, ICL still lags behind LIVE in terms of accuracy.
We also tried to use RICE samples to learn LIVE. However, the results were unsatisfactory, with only 57.68% accuracy on VQAv2, while using random samples to train LIVE got 58.54% accuracy. This may be because using RICE samples will make each training input sequence contain abundant individual characteristics, increasing the difficulty for LIVE to learn common task knowledge, and even causing LIVE to learn spurious correlations between the similar ICDs and the query.